OpenStudy (anonymous):

Solve this Proportion: 1/y = 1/6y-1

OpenStudy (anonymous):

given, 1/y=1/6y-1 or, 1/y-1/6y=-1 or, 5/6y=-1 or, y=-5/6

OpenStudy (anonymous):

1 / y = 1 / 6y - 1 is the proportion.